On July 9, 2019, a truck was purchased for $27,000 with a $2,000 salvage value, is expected to last five years & 40,000 miles. Use the partial year convention to depreciate fractional periods.   Note: you may not require all the given information to solve.                                                                                                                 

The book value of the truck under straight-line at December 31, 2020 after all adjusting entries, will be:

2019...................-250

2020....................-500

27,000-250-500=26,250

2). If the truck was sold on December 31, 2020 for $17,000, compute the gain or loss on the sale.

2020 book value 26,250-17,000=

-9250
Answer #1

Cost price of truck = $27,000

Salvage value = $2,000

Useful life = 5 years

Annual depreciation = (Cost price - Salvage value)/Useful life

= (27,000 - 2,000)/5

= 25,000/5

= $5,000

For the year 2019, depreciation will be charged for 6 months.

Depreciation expense for 2019 = 5,000 x 6/12

= $2,500

The book value of the truck under straight-line at December 31, 2020 after all adjusting entries, will be = Cost price - Accumulated depreciation

= 27,000 - (2,500 + 5,000)

= 27,000 - 7,500

= $19,500

Truck was sold on December 31, 2020 for $17,000

Loss on sale of truck = Book value of the truck at December 31, 2020 - Sale price of truck

= 19,500 - 17,000

= $2,500

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.